### v1

# Mitigating gain calibration errors from EoR observations with SKA1-Low AA*

Abstract
The observation of the redshifted 21-cm signal from neutral hydrogen is a promising probe for understanding the phase transition of the early universe and its evolution during the Cosmic Dawn and the Epoch of Reionisation (EoR). One of the primary obstacles to the statistical detection of the target signal is the presence of residual foreground, which arises from gain calibration errors. Previous studies have shown that gain calibration errors as small as 0.01% can lead to a biased interpretation of the observed signal power spectrum estimation, by nearly an order of magnitude. A recent study further highlights that to retrieve astrophysical parameters accurately, the threshold gain calibration error should be lower than 0.01%. In this work, we develop a post-calibration mitigation strategy that combines foreground subtraction techniques with foreground avoidance to address residual contamination arising from gain calibration errors. We use an end-to-end pipeline 21cmE2E to simulate a realistic sky model and telescope configuration within the 138-146 MHz frequency range. To assess the impact of residual antenna-based gain calibration errors, we perform a detailed power spectrum analysis across several threshold levels. Our analysis shows that the extraction of the target signal over the wavenumber range $0.05\leq k\leq 0.5$ Mpc-1 is possible with a threshold gain calibration error of 1%, although with a significant SNR tradeoff on larger scales. Moreover, this work also offers a comparative assessment of foreground removal and avoidance techniques in the context of future SKA1-Low AA* observations.
1 Introduction
The evolution of the Universe from the post-recombination era to the period following the reionisation of the neutral intergalactic medium (IGM) remains largely unconstrained by direct observations. Theoretical models suggest that the first cosmic structures, comprising the earliest stars and galaxies, formed during this interval, corresponding to redshifts between 30 and 12. The emergence of the first stars, a period referred to as the Cosmic Dawn (CD), was driven by small-scale matter density fluctuations induced by gravitational instabilities. Ionising radiation from these primordial objects gradually ionised the surrounding IGM, initiating the Universe’s final major phase transition. This transformative epoch, known as the Epoch of Reionization (EoR), is believed to span redshifts from approximately 15 to 5 [1, 2].
The neutral hydrogen (Hi) signal from the EoR is expected to be observed at redshifted frequencies in the 50–200 MHz range. Currently, two distinct types of radio telescopes are employed to detect this signal: global signal experiments and interferometers. Global signal experiments use single-radiometer antennas designed to measure the sky-averaged brightness temperature across a broad frequency range. Although EDGES reported an absorption feature centred at 78 MHz [3], subsequent analyses, such as those by SARAS [4], suggest that the best-fit signal may be of non-astrophysical origin. Other single-antenna experiments, including REACH [5] and RHINO [6], are actively pursuing detections of the global signal but are yet to report confirmed results. In parallel, interferometric arrays such as LOFAR [^2] Low Frequency Array, http://www.lofar.org [7], MWA [^3] Murchison Widefield Array, http://www.mwatelescope.org [8], HERA [^4] Hydrogen Epoch of Reionization Array, http://reionization.org [9], uGMRT [^5] Giant Metrewave Radio Telescope , https://www.gmrt.ncra.tifr.res.in/ [10], and the forthcoming SKA [^6] Square Kilometre Array, https://www.skao.int/en [11] aim to detect spatial fluctuations in differential brightness temperature statistically. Although these experiments have placed upper limits on the EoR power spectrum [12, 13, 14, 15], a definitive detection of the Hi signal remains elusive.
The 21-cm signal is susceptible to contamination by bright astrophysical foregrounds, primarily galactic diffuse synchrotron and free-free emissions, along with extragalactic radio sources such as star-forming galaxies and Active Galactic Nuclei (AGN) [16, 17, 18, 19, 20]. These foregrounds are expected to be 4–5 orders of magnitude brighter than the Hi signal. Mitigation of these foregrounds is particularly challenging for interferometric experiments, as global signal experiments benefit from larger beam areas that suppress contributions from bright extragalactic point sources [21]. However, foregrounds in both approaches typically follow a power-law spectrum, leading to inherent spectral smoothness. Since the Hi signal arises against the cosmic microwave background and propagates through an isotropically expanding Universe, its spherical symmetry is preserved in Fourier space, allowing the construction of isotropic image cubes [22]. Unlike foregrounds, the Hi signal exhibits discrete frequency structures due to redshift–frequency mapping, corresponding to line-of-sight (LoS) distances. These contrasting spectral properties form the basis for foreground mitigation strategies.
The effectiveness of foreground mitigation is typically evaluated using cylindrically averaged power spectra (2D PS), which distinctly separate the EoR signal, astrophysical foregrounds, and instrumental systematics in two-dimensional ($k_{\perp},k_{\parallel}$) $k$-space. Smooth foregrounds are expected to dominate low $k_{\parallel}$ modes, contributing mostly around zero delay, while the Hi signal is confined to higher $k_{\parallel}$ values, forming the so-called “EoR window” [23]. The limits in $k_{\perp}$ and $k_{\parallel}$ are governed by the baseline distribution of the array and spectral bandwidth/resolution, respectively. However, instrument chromaticity leads to mode mixing, where foreground power leaks into the EoR window, forming the “foreground wedge” [24, 25, 26]. The statistical detection of the Hi signal is performed using the spherically averaged power spectrum (1D PS).
Foreground mitigation approaches are broadly classified into foreground removal and foreground avoidance [24, 27]. Although both of these approaches have contributed to current upper limits, each has tradeoffs, and no consensus exists on a preferred approach for SKA-Low. Foreground avoidance is robust to systematics but sacrifices sensitivity, especially on large scales, as it discards low $k_{\parallel}$ modes [28]. Conversely, non-parametric subtraction techniques [29, 30, 31, 32, 33, 34], which rely on the frequency-frequency covariance of statistically independent components, are susceptible to signal loss due to leakage of the cosmological signal into the foreground model. The situation is worsened by systematics such as calibration errors, incomplete sky models [35, 36, 37, 38, 39, 40, 41], beam uncertainties [42, 43, 44], mutual coupling [45, 46, 47], and ionospheric distortions [48, 49]. These effects can introduce excess variance or mimic the spectral characteristics of the Hi signal, leading to false detections.
In this work, we focus on mitigating the impact of gain calibration errors using post-calibration residual foreground mitigation techniques. We assess the efficacy of Principal Component Analysis (PCA) and Gaussian Process Regression (GPR) in recovering the 21 cm power spectrum under varying gain error conditions. In addition to purely foreground avoidance or subtraction, inspired by ’hybrid’ foreground mitigation techniques that suggest avoidance along with subtraction [50, 51, 52], we apply PCA and GPR to suppress foreground power at low $k_{\parallel}$ sufficiently to open the EoR window for avoidance-based analysis. We demonstrate that this hybrid approach allows partial recovery of large-scale $k$-modes in the 2D PS, where the Hi signal is strongest, and hence improves the signal-to-noise ratio (SNR). Our primary objective is to mitigate the residual foreground resulting from gain calibration errors as much as possible, while maintaining power spectrum sensitivity.
This paper is structured as follows: Section 2 describes the simulation setup, including both ideal and gain-error–induced SKA1-Low AA* observations. Section 3 outlines the residual foreground mitigation techniques and their implementation. In Section 4, we present the results of the 1D and 2D PS under various calibration conditions. Finally, Section 5 discusses our conclusions and outlines future directions. Throughout the work, we used the best-fitting cosmological parameters from the Planck 2018 results [53].
2 Simulations
This section discusses the simulation framework for introducing residual gain calibration errors into SKA-Low AA* data through synthetic radio observations using end-to-end simulations.
2.1 Synthetic observations
We used a 21-cm end-to-end (21cmE2E) pipeline [39, 48, 54] to perform synthetic observations. This pipeline employs the OSKAR [^7] https://github.com/OxfordSKA/OSKAR software package [55] for simulating visibilities for SKA1-Low configurations. In this simulated observation, the sky was observed with a phase-center at $\alpha$ = 15h00m00s and $\delta$ = -30 for a duration of 4 hours ($\pm$ 2 HA). The observing bandwidth of the lightcone spans 8 MHz with a channel separation of 125 kHz. The observation parameters are listed in the table 1. In the subsection, we describe the realistic sky models and telescope configuration to perform the simulation using the 21cmE2E pipeline.
Parameters Value
[…]
Bandwidth (MHz) 8
Number of Channels 64
Field of View 3
Number of stations 231
Maximum Baseline (km) 3.15
Duration of observation (hrs) 4
Time resolution (sec) 120
Table 1: Observation parameters used for SKA1-Low AA* simulations.
2.1.1 EoR Signal
For Hi simulations, we use the semi-numerical EoR modelling tool, 21cmFAST[56, 57], to generate the brightness temperature map ($\delta T_{b}$). We consider a spatial comoving volume of $500\times 500\,$h-1Mpc2 and a redshift range of 7.4 to 8.2 with pixel size of $232\times 232$ and channel width of 125 kHz. This work used the existing simulated Hi 21-cmmaps from Tripathi et al. [58]. The left plot of Figure 2 shows a 3x3 slice of Hi map at 144.25 MHz. The top-left plot of Figure 3 shows the 2D PS, and the red markers in Figure 4 show the 1D PS of one such realisation of the Hi cube as observed with our telescope model. Since we do not include thermal noise in our data, we consider this to be the ground truth for all power spectrum results.
2.1.2 Point source model
The foreground model used in this study consists exclusively of compact sources extracted from the Tiered Radio Extragalactic Continuum Simulation (T-RECS) catalogue [59]. The sources were redistributed around a declination of -30  and the flux densities were interpolated to 142 MHz by assuming a smooth power law relation with spectral index $\alpha$ equal to -0.8. Our sky model consists of 2522 point sources within our simulated FoV of 4. We refer to Section 2 of [48] for a comprehensive overview of the sky model. It is important to note that the diffuse foreground models are not included in our simulation.
2.1.3 Telescope model
Figure 1: Left : The SKA1-Low AA* stations layout for central 2000 m cutout. Right: UV coverage for the observational parameters in the table 1, with integration time of 4 hours ($\pm 2$ HA). The orange and blue colours represent the (U, V) and (-U,-V) points, respectively.
The SKA1-Low AA* [60] layout is a defined configuration of the SKA1-Low array, comprising 307 stations with a maximum baseline of approximately 73.4 km [61]. It preserves the essential features of the full array design, with a dense central core, where roughly 50% of the stations are packed within a 1 km diameter, and an extended distribution along three logarithmically spaced spiral arms [62]. This layout balances the sensitivity to surface brightness and the angular resolution, crucial for tomographic imaging and statistical studies of EoR [63]. In our work, we restrict the array to stations within a 2 km radius to enhance the sampling of short baselines and thereby improve sensitivity to the cosmologically relevant large-scale 21-cm fluctuations while reducing computational demands. The telescope configuration of SKA1-Low AA* is shown in Figure 1 and the corresponding baseline coverage in the UV plane for an observation time of 4 hours ($\pm 2$ HA).
Figure 2: Left: Simulated $3^{\circ}\times 3^{\circ}$ slice of the observed Hi signal. Right: Residual map after sky-model subtraction with 1% gain calibration error at 144.25 MHz.
2.1.4 Gain Calibration Errors
Gain calibration errors are defined as imperfections in the amplitude and phase of these complex gain terms, which may arise from various instrumental factors, including temperature fluctuations in electronic components, mutual couplings, and time-dependent variations in antenna response. For the scope of this work, we have excluded calibration bandpass errors, assuming that frequency-dependent effects are perfectly corrected and thus do not contribute to the residual errors in our analysis.
The errors due to inaccurate calibration can introduce spurious features that either mimic or entirely obscure the faint 21-cmcosmological signal [40]. Simulations of frequency-independent calibration errors by Morales et al. [64] suggest a more prominent contamination in the reconstructed (image-based) power spectrum than in the delay spectrum estimates. Therefore, we expect that the contamination in our estimation also depends on the quality of the instrumental Point Spread Function (PSF), as pointed out by [35]. Mazumder et al. [39] showed that for SKA1-Low, gain calibration errors as low as 0.1% can significantly contaminate the EoR window, thereby reducing the signal-to-noise ratio. The primary goal of this work is to mitigate the effects of these calibration errors during post-processing, treating them as perturbations in the gain solutions that arise from an unoptimised calibration algorithm.
We define the complex gain $g_{i}$ for the $i^{\text{th}}$ antenna as:
$g_{i}=(a_{i}+\delta a_{i})\exp\left[-i(\phi_{i}+\delta\phi_{i})\right]$ (2.1)
where $a_{i}$ and $\delta a_{i}$ are the amplitude and its error, and $\phi_{i}$ and $\delta\phi_{i}$ are the phase and its error, respectively.
In an ideal situation, $a_{i}=1$ and $\phi_{i}=0$, so we can simplify Equation 2.1 as:
$g_{i}=(1+\delta a_{i})\exp\left[-i\delta\phi_{i}\right]$ (2.2)
Assuming the amplitude and phase errors are independent and drawn from a normal distribution, we rewrite:
$g_{i}=\left(1+\mathcal{N}(0,\sigma_{\text{err}})\right)\exp\left[-i\mathcal{N}(0,\sigma_{\text{err}})\right]$ (2.3)
For each timestamp and throughout the 4-hour observation, these errors are multiplied by the observed visibilities before being subtracted from the sky model. This simulates the percentage inefficiency of the sky-based calibration algorithm and the residual foreground resulting from imperfect subtraction. In this work, we consider three separate cases of a progressively contaminated reconstructed sky varying $\sigma_{\text{err}}$ to 0.1%, 1%, and 10%, respectively.
2.1.5 Imaging
The synthesised maps were generated from the simulated visibilities from the 21cmE2E pipeline using WSCLEAN [65]. The visibilities were gridded, and each sub-band was imaged independently. We stack each sub-band along the frequency axis, making an image cube ($\theta_{x},\theta_{y},\nu$). We apply Gaussian tapering and W-projection to produce naturally weighted dirty images of 4 FoV with a pixel size of $15\times 15$  arcsec. In figure 2, we show one such slice of the image cube of the observed Hi signal maps via SKA-Low simulations. Although we simulate 4 sky areas around the zenith for all the components of the sky signal, we perform the power spectrum estimation using only 3 from the phase-centre, corresponding to SKA-low’s anticipated primary beam [66]. The maximum and minimum $k_{\perp}$ corresponding to image size and PSF are 0.036 Mpc-1 to 0.63 Mpc-1. We find that the effect of PSF starts dominating around $k_{\perp}$ $\sim$ 0.4 Mpc-1. Out of the spherically averaged $k$-modes probable by our observations 0.05 $\leq$ k $\leq$ 1.57 Mpc-1, we focus on large scales of cosmological interest and perform power spectrum analysis for 0.05 $\leq$ k $\leq$ 0.5 Mpc-1.
2.2 Power spectrum estimation
In this section, we provide an overview of image-based (reconstructed) power spectrum estimation used in this work to evaluate cylindrically and spherically averaged 2D and 1D power spectra, respectively.
Let $T(\nu,\mathbf{r})$ be the observed brightness map, and $\tilde{T}(\tau,\mathbf{u})$ be its normalized discrete Fourier transformation, defined as:
$\tilde{T}(\tau,\mathbf{u})=\frac{10^{-26}c^{2}}{2k_{B}N_{v}\nu^{2}\Omega_{PSF}}\sum_{\nu,\mathbf{r}}T(\nu^{\prime},\mathbf{r^{\prime}})e^{-2\iota\pi(\mathbf{u\cdot r^{\prime}+\nu^{\prime}.\tau})}$ (2.4)
where $N_{v}=N_{x}\times N_{y}\times N_{\nu}$ is the number of voxels in our data cube and corresponding voxel resolution as $\delta_{v}=\delta_{x}\times\delta_{y}\times\delta_{\nu}$, where $\delta_{x}$, $\delta_{y}$ and $\delta_{\nu}$ are spatial and spectral resolution, respectively. Before performing the spectral Fourier transform, we divide the temperature voxels with PSF in the $(\nu,\mathbf{u})$ domain and then apply a Blackman-Harris frequency window taper [67]. Connecting radio interferometry to physical cosmology, we define wavenumber $\mathbf{k}=(k_{\perp},k_{\parallel})$ as:
$k_{\perp}=\frac{2\pi|\mathbf{u}|}{D_{M}(z)},\quad k_{\parallel}\approx\frac{2\pi\mathbf{\tau}H_{0}E(z)}{c(1+z)^{2}}$ (2.5)
where $D_{M}(z)$ and $\Delta D$ are conversion factors from angle and frequency to co-moving distance [68, 69], B is the frequency bandwidth, $H_{0}$ is Hubble constant, $\nu_{\rm 21}$ is the frequency of the 21-cm line, $E(z)$ is the dimensionless Hubble parameter defined as $E(z)=(\Omega_{M}(1+z)^{3}+\Omega_{K}(1+z)^{2}+\Omega_{\Lambda})^{1/2}$ [70], and $z$ is the reference redshift at the centre of the observed bandwidth.
We define the cylindrical averaged 2D PS as:
$P(k_{\perp},k_{\parallel})=\mathbb{V}\langle|T(\tau,\mathbf{u})|^{2}\rangle_{k_{\perp},k_{\parallel}},\quad\mathbb{V}=\frac{N_{v}\delta_{v}D^{2}_{M}(z)\Delta D}{\Omega_{PB}B}$ (2.6)
where $\mathbb{V}$ is co-moving volume, $\delta_{x}$, $\delta_{y}$ and $\delta_{\nu}$ are spatial and spectral resolution, respectively.
To quantify the excess residual power in the 2D PS for different cases, we define the absolute fractional error maps as:
$\Delta{P}_{i}(k_{\perp},k_{\parallel})=\frac{|P_{i}(k_{\perp},k_{\parallel})-P_{21}(k_{\perp},k_{\parallel})|}{P_{21}(k_{\perp},k_{\parallel})}$ (2.7)
where ‘i’ corresponds to gain error, PCA or GPR-filtered power spectrum.
The dimensionless 1D PS for spherically averaged $k$-modes $k=\sqrt{k_{\perp}^{2}+k_{\parallel}^{2}}$ is defined as:
$P(k)=\frac{k^{3}}{2\pi^{2}}(\mathbb{V}\langle|T(\tau,\mathbf{u})|^{2}\rangle_{k}),$ (2.8)
We define the wedge line as:
$k_{\parallel}\leq k_{\perp}\frac{D_{M}(z)E(z)H_{0}}{(1+z)c}\sin{\theta}$ (2.9)
where the equation converts to an expression for the horizon line for $\theta$ = $\pi/2$.
To compare the effects of foreground avoidance and subtraction on power spectrum sensitivity and SNR per $k$-bin, we use 21cmSense [^8] https://github.com/rasg-affiliates/21cmSense [71, 72, 28] to evaluate the theoretical thermal noise for 100 hours of SKA1-Low AA* observations. The 1D thermal noise uncertainty is defined as:
$\Delta^{2}_{\rm th}(k)=\frac{k^{3}D^{2}_{M}(z)\Delta D}{2\pi^{2}}\cdot\frac{\Omega^{2}_{P}}{\Omega_{\rm PP}}\cdot\frac{T^{2}_{\rm sys}}{2t}$ (2.10)
Figure 3: Proceeding clockwise from top-left, we show cylindrically averaged 2-D PS for the true Hi (perfect foreground subtraction), and the reconstructed PS of the residual data with 0.1%, 1%, and 10% gain calibration error. The minimum and maximum baselines were taken to be $\sim 48\lambda$ and $\sim 948\lambda$ and the black and white dotted lines represent the horizon limit and horizon plus 0.1 Mpc-1 buffer, respectively.
[…]
The uncertainty due to sample variance is represented as a $2\sigma$ error:
$\sigma^{2}_{i}=\frac{\Delta^{2}(k\in k_{i})}{\sqrt{N_{i}}}$ (2.11)
where $\Delta^{2}(k\in k_{i})$ denotes the spherically averaged power in the $i^{\mathrm{th}}$ $k$-bin, and $N_{i}$ is the number of independent modes contributing to that bin.
The total uncertainty in each $k$-bin is taken as the sum of the thermal and sample variance contributions from equations 2.10 and 2.11.
We define the SNR in each $k$-bin as:
$\mathrm{SNR}(k)=\frac{\Delta_{\rm HI}^{2}(k)}{\Delta_{\rm N}^{2}(k)},$ (2.12)
where $\Delta_{\rm HI}^{2}(k)$ is the Hi power spectrum and $\Delta_{N}^{2}(k)$ is the total uncertainty (sample variance + 100 hr of observation of instrumental noise).
As can be seen from the 2-D cylindrical power spectrum in Figure 3, residual foreground contamination in the EoR window increases with increasing residual gain calibration error. Specifically, for the calibration inaccuracy of 0.1%, the excess residual power manifests itself as a mode-mixing feature [34] at high $k_{\perp}$ modes. Thus, foreground avoidance about the horizon line appears viable (see Section 4.1). However, for higher corruption levels, the familiar foreground wedge, which includes both intrinsic smooth partitions and mode-mixing features, becomes prominent.
The impact of gain errors is also evident in the spherically averaged 1D PS (Figure 4) as excess power above the true Hi signal. This further confirms the contamination introduced by imperfect calibration.
3 Residual Foreground Mitigation
In this section, we provide a brief overview of the foreground mitigation techniques employed in this work, with a focus on their application to EoR observations.
3.1 Foreground Avoidance
As discussed in Section 2.1.4, gain calibration errors increasingly contaminate the EoR window, necessitating a more conservative approach to foreground avoidance. In particular, applying buffers beyond the geometric horizon may be required to exclude contaminated $k$-modes. However, this comes at the cost of reduced sensitivity, particularly on larger scales, due to the loss of low-$k$ modes critical for statistical detection. To investigate this tradeoff, we examine three foreground avoidance scenarios with varying wedge boundaries, similar to the optimistic and pessimistic cases considered in [28].
In the first scenario, we assume ideal subtraction performance and apply no foreground avoidance, thereby retaining all $k$-modes for power spectrum estimation. In the second scenario, we apply a wedge cut at the horizon limit. In the third, more conservative case, we impose an additional buffer of 0.1 Mpc-1 beyond the horizon line.
Figure 5: SNR plot for various cases, defined as the ratio of 1-D Hi power spectrum $\Delta^{2}_{\text{HI}}(k)$ to associated “noise” $\Delta^{2}_{N}(k)$. The noise is the per $k$-bin combined uncertainty due to thermal noise and sample variance, as indicated by vertical error bars in 1-D power spectra (Figure 4). The brackets “hrzn” and “hrzn + 0.1” (dotted lines) indicate that foreground avoidance was applied about the horizon and horizon plus 0.1 Mpc-1 buffer, respectively. Shaded regions represent total sensitivity ($\Delta^{2}_{\text{th}}(k)+\Delta^{2}_{\text{sv}}(k)$) for no avoidance (dark grey), avoidance about the horizon (medium grey), and the horizon plus buffer (light grey).
Figure 5 illustrates the sensitivity curves and resulting SNR for each case. The second scenario (horizon-only avoidance) yields sensitivities nearly identical to the first, deviating only at small scales where thermal noise begins to dominate over sample variance. The average SNR loss across all $k$-modes is approximately 5%. However, in the third scenario, the aggressive avoidance strategy masks a significant portion of the $k$-space, limiting the usable range to $k>0.1$ Mpc-1 and resulting in a substantial average SNR loss of $\sim$30%.
In practice, real observations are often affected by more severe foreground leakage into the EoR window. Thus, careful tuning of the avoidance boundary is crucial to ensure that the retained $k$-modes are dominated by noise rather than systematics. A typical diagnostic technique is to plot 2D log power spectra from cross-correlations of time-interleaved visibilities, identifying EoR-like regions based on the presence of uniformly distributed positive and negative pixels [73]. Alternatively, methods such as filtering [51] or subtraction techniques (explored in the following sections) can be used to clean foreground-contaminated modes within the wedge, effectively improving SNR by expanding the usable region of the EoR window.
3.2 Principal Component Analysis (PCA)
PCA is one of the simplest blind foreground subtraction techniques in 21-cm cosmology [33]. While PCA is classically a linear dimensionality reduction method, it can be repurposed for foreground removal by exploiting the spectral smoothness and frequency correlations of foregrounds. Assuming a noiseless model, our data consist of:
$\vec{d}(\nu,\vec{u})=\vec{d}_{fg}(\nu,\vec{u})+\vec{d}_{21}(\nu,\vec{u}),$ (3.1)
where $\vec{d}$ is the observed data, $\vec{d}_{fg}$ the foreground component, and $\vec{d}_{21}$ the cosmological signal, all as functions of frequency $\nu$ and baseline vector $\vec{u}$.
The PCA steps are as follows:
  • Mean-subtract the data along the LoS: $\vec{d}(\nu,\vec{u})\rightarrow\vec{d}(\nu,\vec{u})-\langle\vec{d}(\nu,\vec{u})\rangle_{\nu}$
  • Compute the covariance matrix: $\boldsymbol{C}=\frac{\vec{d}(\nu,\vec{u})\cdot\vec{d}^{\dagger}(\nu,\vec{u})}{N-1}$
  • Diagonalise $\boldsymbol{C}$ using SVD: $\boldsymbol{\hat{U}^{T}C\hat{U}}=\boldsymbol{\hat{\Lambda}}$, with $\lambda_{i}>\lambda_{i+1}$ for all $i$
  • Choose $N_{fg}$ principal components that capture most of the foreground variance; form matrix $\boldsymbol{W}$ using the corresponding eigenvectors and project onto the data to form the foreground model: $\vec{f}=\boldsymbol{W}^{\dagger}\vec{d}$
  • Subtract the model: $\vec{r}=\vec{d}-\vec{f}$ and use $\vec{r}$ for power spectrum estimation
We choose $N_{fg}$ such that the signal at the largest scale ($k\sim 0.05$ Mpc-1) is not suppressed. Although PCA effectively removes low-rank foreground modes, it struggles to model non-smooth components introduced by mode-mixing or polarisation leakage. More advanced treatments, such as transfer functions [74], can address signal loss but are not pursued here.
3.3 Gaussian Process Regression (GPR)
A Gaussian Process (GP) is defined as a collection of random variables such that any finite subset follows a joint Gaussian distribution [75]. In GPR, we model the data using an infinite-dimensional distribution, defined by a mean and a covariance (kernel) function.
GPR has emerged as a powerful foreground removal technique, both in simulations [76, 34] and in real observations (LOFAR [77, 78, 14]; HERA [79]). Starting again from Eq. 3.1, we define the total kernel:
$K=K_{21}+K_{FG},$ (3.2)
with $K_{FG}=K_{\text{sky}}+K_{\text{mix}}$ separating smooth foregrounds and instrument-induced mode-mixing. We construct a joint distribution for the observed data and the foreground model:
[…]
The foreground mean and covariance estimates are then:
$E[\vec{d}_{fg}]$ $=K_{fg}(K_{fg}+K_{21})^{-1}\vec{d},$ (3.4)
$\text{cov}[\vec{d}_{fg}]$ $=K_{fg}-K_{fg}(K_{fg}+K_{21})^{-1}K_{fg}.$
Subtracting the predicted mean gives the residual:
$r=\vec{d}-E[\vec{d}_{fg}].$ (3.5)
The covariance term provides the uncertainty in the subtraction, which can be used for bias correction.
In practice, kernel components are inferred from the data by optimising hyperparameters. A popular and flexible choice is the Matérn kernel:
$K_{\text{Matern}}(\nu-\nu_{0})=\sigma^{2}\frac{2^{1-\eta}}{\Gamma(\eta)}\left(\sqrt{2\eta}\frac{|\nu-\nu_{0}|}{\ell}\right)^{\eta}K_{\eta}\left(\sqrt{2\eta}\frac{|\nu-\nu_{0}|}{\ell}\right),$
where $\sigma^{2}$ is the variance, $\ell$ is the correlation lengthscale, and $\eta$ governs smoothness. Common choices include: RBF ($\eta\rightarrow\infty$), MATÉRN52 ($\eta=5/2$), MATÉRN32 ($\eta=3/2$), and exponential ($\eta=1/2$). For this work, we use RBF for sky foregrounds, Mat32/52 for mode-mixing, and exponential for the Hi signal.
3.3.1 Application
We apply PCA and GPR to the residual gridded visibilities after dirty image reconstruction, similar to [34, 77]. The data at this stage are one Fourier transform along the LoS away from the delay space.
We use the GPy [^9] https://sheffieldml.github.io/GPy/ package to construct kernels and maximise the log-marginal likelihood (LML):
$\log p(\mathbf{d}|\nu,\theta)=-\frac{1}{2}\left(\mathbf{d}^{T}K^{-1}\mathbf{d}+\log|K|+n\log 2\pi\right),$ (3.6)
where $\mathbf{d}$ is the data, $K$ the total kernel, $n$ the number of data points, and $\theta$ the hyperparameters.
Hyperparameter optimisation is performed using the emcee MCMC sampler. The best-fit values are used to construct the foreground model, which is subtracted from the visibilities. We verify the robustness of sampling by exploring different prior ranges.
Initial runs showed that with increasing gain error, the Hi coherence scale hyperparameter $l_{21}$ converged to the prior upper bound. Inspired by [80], we adopt a gamma prior with an expectation value of 0.1 MHz (the known Hi scale), corresponding to $\Gamma(0.5,5)$. In the Table 2, we listed the hyperparameters used in GPR to subtract the residual foreground. The corner plots of the posterior probability distributions of GPR hyperparameters after MCMC sampling for different calibration errors are shown in Figure 6.
Parameters $\eta$ Prior Posterior
1% 10% 1% 10%
$\sigma_{\text{sky}}/\sigma_{n}$ $\infty$ $\mathbb{U}(0.1,5)$ $\mathbb{U}(0.1,1)$ 1.03 0.48
$l_{\text{sky}}$ (MHz) $\mathbb{U}(5,50)$ $\mathbb{U}(1,25)$ 24.11 3.03
$\sigma_{21}/\sigma_{n}$ 1/2 $\mathbb{U}(10^{-6},10^{-1})$ $\mathbb{U}(10^{-8},10^{-1})$ $10^{-3}$ $10^{-7}$
$l_{21}$ (MHz) $\Gamma(0.5,5)$ $\Gamma(0.5,5)$ 26.39 11.53
Table 2: The GPR kernel priors and posteriors for 1% and 10% gain calibration error cases.
Figure 6: Posterior distributions of GPR hyperparameters after MCMC sampling for 1% (left) and 10%(right) gain calibration errors. Contours represent confidence levels of 11.8%, 39.3%, 67.5%, and 87.4%.
4 Results
In this section, we discuss the results obtained after application of three distinct residual foreground mitigation strategies: avoidance, subtraction, and a hybrid approach. The subtraction methods use statistical modelling of foregrounds using either PCA or GPR, while the avoidance technique excludes $k$-modes below a designated wedge boundary, thereby separating regions heavily contaminated by residual foregrounds. The hybrid approach sequentially combines both subtraction and avoidance strategies. We aim to evaluate the robustness of each technique based on its effectiveness in recovering the true underlying Hi signal. In the following subsections, we discuss how different foreground mitigation strategies recover the target signal.
4.1 Case I: Gain Error of $\sim 0.1\%$
In this case, we incorporate a gain calibration error of 0.1% and corrupt the simulated observed visibilities accordingly. The impact of a 0.1% gain calibration error on 1D PS is presented in Figure 4. In addition, the absolute fractional error of 2D PS, computed using Equation 2.7, is shown in the right panel of Figure 7. The 1D PS indicates that on larger scales ($k\leq 0.2$ Mpc-1), the residual power closely matches the target signal level, even without applying any mitigation. However, on smaller scales ($k>0.2$ Mpc-1), the residual power underestimates the true signal by nearly an order of magnitude. By applying foreground avoidance techniques using the horizon line, as shown in the left panel of Figure 7, the signal recovery improves significantly throughout the $0.05\leq k\leq 0.5$ Mpc-1 range, despite an estimated loss of $\sim$5% in SNR (see Section 3.1). For this specific scenario, neither subtraction nor hybrid mitigation methods were employed.
Figure 7: Power spectrum analysis after residual foreground mitigation for 0.1% gain calibration error. Left: Absolute fractional error map $\Delta P_{0.1\%}$ for sky-model subtracted visibilities, evaluated according to equation 2.7. The horizon limit (dotted black line) is used as the wedge boundary, with shaded regions indicating masked $k$-modes. Right: The square markers show the 1D PS before and after foreground avoidance. The grey-shaded region represents the true HI signal. Binning and error bars follow Figure 4.
4.2 Case II: Gain Error of $\sim 1\%$
Figure 8: The 2D absolute fractional error maps $\Delta P_{i}(k_{\perp},k_{\parallel})$ for 1% (top row) and 10% (bottom row) gain calibration errors. Columns (from left to right) correspond to sky-model subtraction, PCA subtraction, and GPR subtraction. The dotted black lines indicate the wedge boundaries used for avoidance; the shaded areas represent the masked $k$-modes for that particular case. Bluer regions indicate better agreement with the true Hi signal. Corresponding 1D PS are shown in Figures 9 and 10.
To further quantify the impact of calibration systematics, we introduce a gain calibration error of 1% and corrupt the simulated visibilities accordingly. The resulting cylindrical power spectrum is shown in the bottom-right panel of Figure 3, while the corresponding spherically averaged power spectrum is presented in Figure 4. A comparison between the residual power spectrum, affected by a 1% gain calibration error, and the true Hi power spectrum reveals substantial discrepancies across all $k$ modes, with residual power consistently exceeding the true signal by a significant margin. Relative to the case of 0.1% error, these deviations are substantially larger, clearly demonstrating that foreground avoidance alone is insufficient for reliable signal recovery under 1% gain calibration errors.
The plots in the top row in Figure 8 show the variation in 2D absolute fractional error maps for 1% calibration gain error, after PCA subtraction, and after GPR subtraction, respectively. One can see that the subtraction techniques effectively reduce the absolute fractional error by up to an order of magnitude in the low $k_{\perp}$ range ($\leq 0.1$Mpc-1). Although GPR performs marginally better for higher $k_{\perp}$ modes than PCA, the mode-mixing component dominates, and we do not observe a significant improvement with either method. The corresponding spherically averaged power spectrum after the subtraction techniques is presented in the left plot of Figure 9.
Using avoidance around the horizon line or a 0.05 Mpc-1 buffer, our model fails to recover the signal on large scales. Avoidance with a buffer of 0.1 Mpc-1 (Figure 8, top-left) successfully recovers the signal but results in an average SNR loss of at least 30% (see Section 3.1), as indicated by the cyan markers in the right plot of Figure 9.
Lastly, using a hybrid technique of applying avoidance (with a 0.05 Mpc-1 buffer) to PCA/GPR residuals, it enables us to successfully recover the true signal Hiin the available $k$ modes, as represented by the green and blue markers in the right plot of Figure 9.
Figure 9: Spherically averaged 1D PS after residual foreground mitigation for 1% gain calibration error. Left (Subtraction only): Star markers denote PCA and GPR-subtracted residuals. The shaded region represents the true Hi signal. Right (Hybrid): The cyan markers show avoidance using a 0.1 Mpc-1 buffer; green and blue markers represent hybrid (subtraction + 0.05 Mpc-1 buffer avoidance) cases. The orange and brown crossed-out regions represent the $k$-modes that are inaccessible due to avoidance with 0.05 and 0.1 Mpc-1 buffers, respectively. Markers for GPR are offset by 5% in $k$ for visibility.
4.3 Case III: Gain Error of $\sim 10\%$
In this case, we corrupted the observed visibilities with a 10% gain calibration error to assess its impact on the recovered power spectrum. The effects are illustrated in the cylindrical power spectrum (bottom-left panel of Figure 3) and the spherically averaged power spectrum (Figure 4). We further examined the effectiveness of subtraction-based mitigation techniques in this high-error scenario. The plots in the bottom row in Figure 8 show the variation in the 2D absolute fractional error maps for 10% calibration gain error and after using subtraction methods. Applying GPR and PCA independently shows that subtraction alone is insufficient. GPR severely overestimates the signal, with strong suppression on larger scales, while PCA fails to account for the strong mode-mixing component on smaller scales. Hence, neither method adequately captures the foreground covariance at this level of contamination. Foreground avoidance with a 0.1 Mpc-1 buffer also fails to recover the signal for $k\leq 0.2$ Mpc-1. However, 1D PS results as shown in Figure  10 combining avoidance with a 0.1 Mpc-1 buffer along with GPR subtraction, though still inaccessible for $k<0.1$ Mpc-1, perform better on smaller scales compared to hybrid PCA ($n_{\text{fg}}=3$).
Figure 10: Spherically averaged 1D PS after residual foreground mitigation for 10% gain calibration error. Hybrid: The cyan markers indicate avoidance using a 0.1 Mpc-1 buffer. Blue and green markers show hybrid PCA and GPR results. The grey-shaded region represents the true HI signal while the brown, crossed-out regions represent the inaccessible $k$-modes due to avoidance with a 0.1 Mpc-1 buffer.
As can be seen in Figure 11, the frequency-frequency covariance map of the foreground model retrieved after GPR exhibits smoothly varying features across the bandwidth, reminiscent of foregrounds, for a 1% gain calibration error, as opposed to comparatively less smooth features for a 10% error. For both contamination levels, incorporating additional mode-mixing kernels (e.g., MAT-3/2 and 5/2) into the GPR model further suppressed large-scale power on smaller scales, while improving the Bayesian evidence and recovering $l_{21}$ more accurately.
Figure 11: Frequency-frequency covariance of the GPR-reconstructed residual foregrounds for 1% (left) and 10% (right) gain errors. Spectral smoothness degrades at higher gain errors, evident from diagonally auto-correlated structures in the 10% case.
5 Summary and Conclusion
In this study, we examined the impact of such gain errors on statistical detection of the 21-cm signal and assessed mitigation strategies based on residual foreground techniques. We used the 21cmE2E pipeline to simulate the mock observation data for SKA1-Low AA* configuration. We computed 2D and 1D PS from image cubes following three mitigation strategies: foreground subtraction (PCA, GPR), foreground avoidance, and a hybrid combination. Key findings for different gain error levels are summarised below:
  • Gain Error $\sim 0.1$%: Foreground avoidance about the horizon line enables successful recovery of the Hi signal across $k$-modes, with only $\sim$5% SNR loss. Subtraction or hybrid methods were unnecessary in this low-contamination regime.
  • Gain Error $\sim 1$%: GPR outperforms PCA on larger scales ($k\leq 0.1$ Mpc-1), but both methods fall short on smaller scales. Foreground avoidance with a 0.1 Mpc-1 buffer, or hybrid mitigation with a 0.05 Mpc-1 buffer post-subtraction, enables signal recovery across accessible $k$-modes, albeit with an average SNR loss of approximately 30%.
  • Gain Error $\sim 10$%: Subtraction (PCA or GPR) and avoidance (0.1 Mpc-1 buffer) alone fail to recover the signal. GPR especially overestimates small-scale power while suppressing large-scale modes. However, hybrid mitigation using GPR combined with avoidance successfully recovers the signal in the range $0.1\leq k\leq 0.5$ Mpc-1. Again, this recovery comes at an average 30% loss of SNR in the accessible $k$-modes.
Our results show that increasing gain errors contaminate the EoR window and erode the spectral smoothness of foregrounds, thereby limiting the effectiveness of standard subtraction-based methods. This raises a key question for future investigations: to what extent do foregrounds retain their smoothness when subjected to realistic instrumental effects and calibration systematics? Although the standard GPR technique [34] has the capability to model non-smooth foreground components, kernel selection remains a challenging task, especially in high-contamination scenarios [81]. Our findings suggest that a hybrid approach, combining model-based subtraction with strategic avoidance, can recover the signal Hi even under moderate to severe contamination. However, this comes at the cost of reduced power spectrum sensitivity and signal-to-noise ratio on cosmologically relevant scales. Therefore, careful inspection of the wedge boundary and the spectral windowing becomes essential [82].
Future work will include extending this framework to account for other observational systematics such as thermal noise, source position offsets, primary beam uncertainties, and bandpass calibration errors. We also plan to move toward Bayesian parameter estimation using emulator-based inference pipelines (e.g., [58]), and apply our methodology to real datasets, including upcoming Band-2 (120-250 MHz) uGMRT observations at $z=5.3,6.3,\,9.0\,\text{and}\,10.2$ (Sagar et al., in preparation).
Acknowledgments
EB and AT thank the Indian Institute of Technology Indore for providing funding for this study in the form of a Teaching Assistantship. SKP acknowledges the financial support from the Department of Science and Technology, Government of India, through the INSPIRE Fellowship [IF200312]. The authors acknowledge the use of facilities procured through the funding via the Department of Science and Technology, Government of India, sponsored DST-FIST grant no. SR/FST/PSII/2021/162 (C) awarded to DAASE, IIT Indore.
Software: We would like to acknowledge the usage of PYTHON programming language ( https://www.python.org/ ), numpy ([83]), scipy ([84]), astropy ([85]), emcee ([86]), and matplotlib ([87]) in this work.